My Square has an area of 122 Meters Squared. The Length is 6 but i cant find the width. i tried dividing 122 by 6 but it didnt help.
Answer by Alan3354(69443)

My Square has an area of 122 Meters Squared. The Length is 6 but i cant find the width. i tried dividing 122 by 6 but it didnt help.
--------------
Area = L*W = 122
6W = 122
W = 122/6 = 61/3 meters = 20.333... meters
===========================
That's the width, whether it helps or not.